# Modeling Ultra-High-Energy Cosmic Rays propagation using the input from Configuration Interaction Shell Model

### v1

@@ Methods
We employ the Configuration Interaction Shell Model (CI-SM) approach to provide predictions of $E1$ dipole response for $p$ and sdsd-shell nuclei, with mass number AA between 7 and 40. Theoretical predictions are compared to available data and to existing predictions from phenomenological and microscopic models. Finally, the impact of using of CI-SM PSF on the predicted propagation of a 40Ca UHECR source is studied.
@@ Conclusions
The CI-SM method is shown to be a valuable alternative to provide PSF for the modeling of the propagation of the UHECR, so far only achieved with linear-response models and phenomenological models of nuclear strength functions. While this work covers a considerable fraction of nuclei necessary in such applications (namely $p$ and sdsd-shell nuclei), future work will also treat pfpf-shell systems and hence complete the CI-SM systematics of dipole PSF necessary for such applications.
@@ I Introduction
The average response of a nucleus to an electromagnetic perturbation is characterized by its photon strength function (PSF). PSFs are key quantities in many areas of nuclear physics, serving as essential tools for probing nuclear structure, from single-particle motion to collective excitations, as well as in the description of astrophysical processes. Among all types and multipolarities of electromagnetic response, the electric dipole (E1) response is usually the most probable, especially in the $\gamma$-ray region of 10-30 MeV encompassing the well-known giant dipole resonance which is a fully collective dipole mode with protons and neutrons oscillating out of phase [1]. In neutron-rich nuclei, the oscillation of the excess neutrons against the isospin-saturated core is predicted to generate a low-lying E1 strength, referred to as the pygmy dipole resonance (PDR) [2]. Studying the PDR provides a means to probe the neutron-skin thickness in medium to heavy nuclei [3, 4], places constraints on the nuclear symmetry energy [5], and provides insight into the properties of neutron stars [6]. PSFs are also key ingredients to the calculations of radiative neutron capture cross sections, which are needed e.g. to model the intermediate [7, 8] or rapid [9, 10, 11] neutron capture processes of nucleosynthesis. Last but not least, PSFs are necessary in the simulations of the propagation of ultra-high energy cosmic rays (UHECRs). This topic has recently gained significant attention through the PANDORA project (Photo-Absorption of Nuclei and Decay Observation for Reactions in Astrophysics) [12], which was initiated to study the photo-nuclear reactions of light nuclei both experimentally and theoretically. Part of this initiative focuses on improving the characterization of UHECR’s photo-interactions, a crucial component in the theoretical modeling of their propagation through extragalactic space. UHECR photo-interactions can significantly vary depending on the chosen PSF inputs in the calculations of photo-nuclear cross sections [13, 14]. To investigate the propagation of UHECR, PSFs of nuclei with mass $A\leq 60$ are needed [12]. Experimental photodata are scarce and often mutually contradictory in this target mass region. Therefore, one must resort to theoretical predictions of PSFs. Large scale PSF predictions primarily rely on parameterized models like the simple modified Lorentzian model (SMLO) [15, 16] and on declinations of linear response approaches such as the quasi-particle random phase approximation (QRPA) [17, 18] and its quasi-particle finite amplitude method (QFAM) formulation [19, 20], which has the advantage of being computationally cheaper than the former. Provided that phenomenological adjustments, such as mass dependent centroid shift and folding width, are added, such microscopic predictions are efficient for mid-mass and heavy nuclei, which display mainly smooth bulk-like responses. On the other hand, in the case of light nuclei, linear response fails to reproduce the PSF fragmentation [21]. To have reliable PSFs predictions for light nuclei, one may rely on microscopic approaches that account for correlations beyond the harmonic approximation of the QRPA, such as second QRPA [22], QRPA plus phonon coupling [23], projected generator coordinate method [24, 25], or the Configuration Interaction Shell Model (CI-SM)[26]. The large-scale evaluations of PSF within those models are, however, not yet accessible due to numerical complexity. In particular, calculations of electric dipole excitation strength functions within the CI-SM framework are relatively scarce in the literature, mainly due to the challenges associated with deriving multi-shell valence-space interactions and the large dimensions of the Hamiltonian matrices involved. Recently, the $E1$ PSFs of all long-lived sdsd-shell nuclei have been calculated within the CI-SM framework [21]. Building on these efforts, the present study aims at: 1) expanding the CI-SM investigation through a systematic analysis of the $E1$ photoabsorption strength function in $p$ and sdsd-shell nuclei; 2) benchmarking the results against experimental data; 3) comparing to existing theoretical predictions, and 4) estimating the impact of the CI-SM predictions on the propagation of UHECRs.
@@ II.1 Theoretical framework
Shell-model calculations performed in this work span masses between $A=7$ and $A=40$, which means $p$-shell and sdsd-shell nuclei. This requires two distinct model spaces to achieve the description of $E1$ excitations of these systems. Our approach to derive the $E1$ strength of sdsd-shell nuclei, using the PSDPF interaction [28], was recently presented in detail in Ref. [21]. To compute the $E1$ response of $p$-shell nuclei we used the WBP effective Hamiltonian defined in the in the spsdpfs-p-sd-pf model space [29]. In both model spaces, we allow only for $0\hbar\omega$ ground states (GS) and $1\hbar\omega$ excited states. The reduced transition probability in CI-SM is calculated as
@@
which has been shown to provide the correct behavior at low excitation energy [30, 31, 18]. The width parameter is set to $\Gamma=1$ MeV for sdsd-shell nuclei, with the exception of 40Ca, for which $\Gamma=3.5$ MeV is adopted to account for the bunching of the strength into a few $B_{0\nu}$ peaks, a consequence of the limited correlations due to the complete filling of the sdsd-shell. For $p$-shell nuclei, a width of $\Gamma=3.5$ MeV is employed in all cases. We present the results of the computations as PSFs (in MeV-3) defined by the relation [32]:
@@ II.2 Normalization of the PSF strength
Systematic overestimation of the $E1$ total strength function appears to be a common feature with available shell-model calculations [21, 33, 34, 35, 36, 37]. As in the case of $M1$, $E2$, $E3$, and Gamow-Teller transitions [26, 38, 39, 40, 41], the $E1$ operator must be renormalized, or “dressed,” to incorporate correlations beyond the model space. Microscopic methods for mapping an operator into a model space include Lee–Suzuki approach [42] and similarity renormalization group techniques [43]. Phenomenological effective charges are often favored over these methods due to their complexity and the difficulties associated with managing multi-shell model spaces. In Ref. [21], we determined the $E1$ effective charges for sdsd-shell nuclei as $(e_{\textrm{eff}}^{n},e_{\textrm{eff}}^{p})=(-0.8Z/A,0.8N/A)$, corresponding to an overall reduction of the total strength by a factor of $Q^{2}=0.64$. In this work, we adopt a slightly different value of $Q^{2}$ determined as:
@@
where $S_{1}$ is the energy-weighted sum rule (EWSR) from Eq. 4 obtained using the bare dipole operator, $S_{\textrm{\scriptsize TRK}}=14.8NZ/A\quad e^{2}\textrm{fm}^{2}\penalty 10000\ \textrm{MeV}$ is the Thomas-Reich-Kuhn (TRK) sum rule, and $\kappa$ an enhancement factor [44]. It is known that the enhancement factor critically depends on an energy cutoff $E_{\gamma}^{\textrm{max}}$ [1, 45, 46, 47]. Typically for $E_{\gamma}^{\textrm{max}}\sim 140$ MeV, $\mathcal{O}(\kappa)=1$ while if $E_{\gamma}^{\textrm{max}}\sim E_{\textrm{\scriptsize GDR}}$, $\kappa$ is much smaller. In the present calculations, we set the energy range of interest to $E_{\gamma}^{\textrm{max}}=50$ MeV. A phenomenological adjustment comes in the enhancement factor assumed in the present calculations, as $\kappa=0.14$ for sdsd-shell nuclei [21] and $\kappa=0$ for $p$-shell nuclei.
@@ III.1 $p$-shell nuclei
III.2 sdsd-shell nuclei
Figure 2: Same as Fig. 1 for sdsd-shell nuclei
@@
In Ref. [21], we presented systematic calculations of PSFs in long-lived sdsd-shell nuclei and compared them to available data and D1M+QRPA predictions from Ref. [18]. In this work we extended the CI-SM calculations to provide predictions for all 141 existing isotopes that can be described in this valence space. Figs. 2 and 3 compare our CI-SM results for sdsd-shell nuclei close to the stability line with those obtained with other models, namely D1M+QRPA, SMLO [15, 16] and RQFAMz [49]. It can be observed that, while in certain cases the CI-SM predictions are in reasonable agreement with the SMLO and RQFAMz PSFs, in many others the CI-SM results exhibit a higher degree of fragmentation and slight differences in strength magnitude, width, and centroid, which may influence photodisintegration cross sections. Fig. 3 shows the evolution of the centroid and the width as a function of AA. The decrease of the centroid energy with increasing mass is consistently observed across all three microscopic models with a similar trend. It is worth noting that both D1M+QRPA and RQFAMz include global phenomenological adjustments affecting not only the centroid position but also the folding width. The CI-SM centroids exhibit less smooth variations compared to linear-response-based approaches, as they can incorporate rapid changes in the shell structure and in the GS correlations. Similarly, the CI-SM widths are more scattered and show no clear systematic trend, whereas the D1M+QRPA and RQFAMz widths exhibit a smooth mass dependence. In Ref. [50], we presented a systematic comparison of CI-SM centroids and widths with experimental data. The CI-SM centroid predictions appeared slightly lower than the measured values, with a root-mean-square (rms) deviation of $0.85$ MeV across the 26 cases (see Table 1). For the CI-SM widths, the corresponding rms deviation is $0.50$ MeV, with no systematic under- or over-estimate [^1] Note that the centroid and the width as defined in Eq. 5 are independent of the normalization of the PSF strength described in Sec. II.2.. In CI-SM, we account for the Landau damping as well as most of the spreading width; therefore, the calculated width is almost not sensitive to the folding, which only accounts for the escape width. For comparison, the rms deviation in D1M+QRPA predictions amounts to $1.71$ MeV on the centroids and $0.49$ MeV on the widths, while RQFAMz leads to an rms deviations of $0.89$ MeV on the centroids and $0.48$ MeV on the widths. Hence, even without empirical adjustments to the centroids or widths, the CI-SM predictions remain slightly more accurate than RQFAMz PSFs for sdsd-shell nuclei. One may note that in many cases the experimental PSFs themselves are mutually inconsistent (see Fig. 2), so these rms values should be interpreted with caution. We shall also recall that the shell-model PSDPF effective interaction [28] was only adjusted to reproduce the low-energy spectroscopic data. Therefore, a possible improvement in the PSF predictions could be to incorporate in the fit of the interaction also the measured centroids and widths. From a purely practical perspective one could apply a simple centroid shift, in analogy with the procedure used in QFAM/QRPA calculations. Because the uncorrected CI-SM reproduces the known centroids better than QRPA-based models with empirical corrections, we chose to base our predictions for nuclei used in Sec. IV on the bare CI-SM results.
@@
Figure 3: Panel (a): centroids (up to $30$ MeV) of $E1$ distributions as a function of the mass number AA for long-lived sdsd-shell nuclei. Panel (b): width (up to $30$ MeV) of $E1$ PSFs as a function of the mass number AA for long-lived sdsd-shell nuclei. In red CI-SM, in orange D1M+QRPA [31], in blue RQFAMz results [49].
@@ III.3 O and Ne isotopic chains
After comparing nuclei close to the stability line, we turn to the predictions in the neutron-rich region, where nuclear models tend to show the largest discrepancies. In Figs. 4 and 5, we compare the PSF predictions from all four models for oxygen and neon isotopic chains. As observed earlier, the CI-SM PSFs exhibit significantly more fragmentation than the linear-response-based approaches. While D1M+QRPA and RQFAMz show a smooth evolution of the PSFs with increasing neutron number, the CI-SM dipole responses display sharper variations. Clearly, the SMLO predictions were not tailored for neutron-rich nuclei and therefore cannot account for any PDR. In all three microscopic models, we observe an accumulation of low-lying dipole strength with increasing neutron number, characteristic of the emergence of a PDR. Although it may seem at first sight that no PDR appears in the D1M+QRPA case, it is, in fact, dissolved in the GDR tail. In Refs. [51, 50], we showed that the CI-SM prediction for the PDR in 26Ne is very close to the experimental values (both strength and centroid), which is promising for extending such predictions to other neutron-rich nuclei. Fig. 6 shows the centroids (calculated up to $30$ MeV) for all four models. As can be seen, D1M+QRPA, RQFAMz, and SMLO exhibit the same trend as a function of neutron excess, displaying a gradual decrease. The difference in absolute values between D1M+QRPA and RQFAMz can be attributed to the fact that RQFAMz predicts significantly more strength at high energies (20203030 MeV) than D1M+QRPA. Additionally, phenomenological corrections made in D1M+QRPA PSFs optimize the centroid energies in medium and heavy mass nuclei but tend to underestimate them for light nuclei [18]. Interestingly, the CI-SM centroids follow a very different trend from the other models, exhibiting a much steeper decrease with increasing AA. This effect is also evident in Fig. 7, where we show the fraction of the PDR EWSR calculated up to $15$ MeV relative to the TRK sum rule [^2] The choice of $15$ MeV as the upper bound, rather than $10$ MeV, is motivated by the fact that the experimental values in Ref. [52] are reported up to this energy. As a consequence, this strength is not purely PDR since it incorporates part of the GDR tail.. Clearly, our calculations predict a stronger increase in the low-lying strength as a function of neutron excess. For the oxygen isotopic chain, experimental data on the PDR are available for a few isotopes [52]. As can be seen in the left panel of Fig. 7, the CI-SM predictions agree quite well with the experiment up to 20O. While they deviate for 21O and 22O, the overall trend remains consistent with the data. This deviation can be attributed to the systematic underestimation of the PSF centroid discussed in the previous section. Indeed, introducing a 11 MeV upward shift would lead to a much better agreement with the data. While the differences in nuclear models observed for neutron-rich nuclei appear substantial, they have no impact on the modeling of the UHECR, which is dominated by the description of stable and proton-rich nuclei (see Sec. IV). Similar model-to-model variations also appear on the proton-rich side, although the proximity of the proton drip line suppresses the most divergent cases.

### v2

@@ Methods
We employ the Configuration Interaction Shell Model (CI-SM) approach to provide predictions of $E1$ dipole response for $p$ and $sd$-shell nuclei, with mass number $A$ between 7 and 40. Theoretical predictions are compared to available data and to existing predictions from phenomenological and microscopic models. Finally, the impact of using of CI-SM PSF on the predicted propagation of a 40Ca UHECR source is studied.
@@ Conclusions
The CI-SM method is shown to be a valuable alternative to provide PSF for the modeling of the propagation of the UHECR, so far only achieved with linear-response models and phenomenological models of nuclear strength functions. While this work covers a considerable fraction of nuclei necessary in such applications (namely $p$ and $sd$-shell nuclei), future work will also treat $pf$-shell systems and hence complete the CI-SM systematics of dipole PSF necessary for such applications.
@@ I Introduction
The average response of a nucleus to an electromagnetic perturbation is characterized by its photon strength function (PSF). PSFs are key quantities in many areas of nuclear physics, serving as essential tools for probing nuclear structure, from single-particle motion to collective excitations, as well as in the description of astrophysical processes. Among all types and multipolarities of electromagnetic response, the electric dipole (E1) response is usually the most probable, especially in the $\gamma$-ray region of 10-30 MeV encompassing the well-known giant dipole resonance which is a fully collective dipole mode with protons and neutrons oscillating out of phase [1]. In neutron-rich nuclei, the oscillation of the excess neutrons against the isospin-saturated core is predicted to generate a low-lying E1 strength, referred to as the pygmy dipole resonance (PDR) [2]. Studying the PDR provides a means to probe the neutron-skin thickness in medium to heavy nuclei [3, 4], places constraints on the nuclear symmetry energy [5], and provides insight into the properties of neutron stars [6]. PSFs are also key ingredients to the calculations of radiative neutron capture cross sections, which are needed e.g. to model the intermediate [7, 8] or rapid [9, 10, 11] neutron capture processes of nucleosynthesis. Last but not least, PSFs are necessary in the simulations of the propagation of ultra-high energy cosmic rays (UHECRs). This topic has recently gained significant attention through the PANDORA project (Photo-Absorption of Nuclei and Decay Observation for Reactions in Astrophysics) [12], which was initiated to study the photo-nuclear reactions of light nuclei both experimentally and theoretically. Part of this initiative focuses on improving the characterization of UHECR’s photo-interactions, a crucial component in the theoretical modeling of their propagation through extragalactic space. UHECR photo-interactions can significantly vary depending on the chosen PSF inputs in the calculations of photo-nuclear cross sections [13, 14]. To investigate the propagation of UHECR, PSFs of nuclei with mass $A\leq 60$ are needed [12]. Experimental photodata are scarce and often mutually contradictory in this target mass region. Therefore, one must resort to theoretical predictions of PSFs. Large scale PSF predictions primarily rely on parameterized models like the simple modified Lorentzian model (SMLO) [15, 16] and on declinations of linear response approaches such as the quasi-particle random phase approximation (QRPA) [17, 18] and its quasi-particle finite amplitude method (QFAM) formulation [19, 20], which has the advantage of being computationally cheaper than the former. Provided that phenomenological adjustments, such as mass dependent centroid shift and folding width, are added, such microscopic predictions are efficient for mid-mass and heavy nuclei, which display mainly smooth bulk-like responses. On the other hand, in the case of light nuclei, linear response fails to reproduce the PSF fragmentation [21]. To have reliable PSFs predictions for light nuclei, one may rely on microscopic approaches that account for correlations beyond the harmonic approximation of the QRPA, such as second QRPA [22], QRPA plus phonon coupling [23], projected generator coordinate method [24, 25], or the Configuration Interaction Shell Model (CI-SM)[26]. The large-scale evaluations of PSF within those models are, however, not yet accessible due to numerical complexity. In particular, calculations of electric dipole excitation strength functions within the CI-SM framework are relatively scarce in the literature, mainly due to the challenges associated with deriving multi-shell valence-space interactions and the large dimensions of the Hamiltonian matrices involved. Recently, the $E1$ PSFs of all long-lived $sd$-shell nuclei have been calculated within the CI-SM framework [21]. Building on these efforts, the present study aims at: 1) expanding the CI-SM investigation through a systematic analysis of the $E1$ photoabsorption strength function in $p$ and $sd$-shell nuclei; 2) benchmarking the results against experimental data; 3) comparing to existing theoretical predictions, and 4) estimating the impact of the CI-SM predictions on the propagation of UHECRs.
@@ II.1 Theoretical framework
Shell-model calculations performed in this work span masses between $A=7$ and $A=40$, which means $p$-shell and $sd$-shell nuclei. This requires two distinct model spaces to achieve the description of $E1$ excitations of these systems. Our approach to derive the $E1$ strength of $sd$-shell nuclei, using the PSDPF interaction [28], was recently presented in detail in Ref. [21]. To compute the $E1$ response of $p$-shell nuclei we used the WBP effective Hamiltonian defined in the in the spsdpfs-p-sd-pf model space [29]. In both model spaces, we allow only for $0\hbar\omega$ ground states (GS) and $1\hbar\omega$ excited states. The reduced transition probability in CI-SM is calculated as
@@
which has been shown to provide the correct behavior at low excitation energy [30, 31, 18]. The width parameter is set to $\Gamma=1$ MeV for $sd$-shell nuclei, with the exception of 40Ca, for which $\Gamma=3.5$ MeV is adopted to account for the bunching of the strength into a few $B_{0\nu}$ peaks, a consequence of the limited correlations due to the complete filling of the $sd$-shell. For $p$-shell nuclei, a width of $\Gamma=3.5$ MeV is employed in all cases. We present the results of the computations as PSFs (in MeV-3) defined by the relation [32]:
@@ II.2 Normalization of the PSF strength
Systematic overestimation of the $E1$ total strength function appears to be a common feature with available shell-model calculations [21, 33, 34, 35, 36, 37]. As in the case of $M1$, $E2$, $E3$, and Gamow-Teller transitions [26, 38, 39, 40, 41], the $E1$ operator must be renormalized, or “dressed,” to incorporate correlations beyond the model space. Microscopic methods for mapping an operator into a model space include Lee–Suzuki approach [42] and similarity renormalization group techniques [43]. Phenomenological effective charges are often favored over these methods due to their complexity and the difficulties associated with managing multi-shell model spaces. In Ref. [21], we determined the $E1$ effective charges for $sd$-shell nuclei as $(e_{\textrm{eff}}^{n},e_{\textrm{eff}}^{p})=(-0.8Z/A,0.8N/A)$, corresponding to an overall reduction of the total strength by a factor of $Q^{2}=0.64$. In this work, we adopt a slightly different value of $Q^{2}$ determined as:
@@
where $S_{1}$ is the energy-weighted sum rule (EWSR) from Eq. 4 obtained using the bare dipole operator, $S_{\textrm{\scriptsize TRK}}=14.8NZ/A\quad e^{2}\textrm{fm}^{2}\penalty 10000\ \textrm{MeV}$ is the Thomas-Reich-Kuhn (TRK) sum rule, and $\kappa$ an enhancement factor [44]. It is known that the enhancement factor critically depends on an energy cutoff $E_{\gamma}^{\textrm{max}}$ [1, 45, 46, 47]. Typically for $E_{\gamma}^{\textrm{max}}\sim 140$ MeV, $\mathcal{O}(\kappa)=1$ while if $E_{\gamma}^{\textrm{max}}\sim E_{\textrm{\scriptsize GDR}}$, $\kappa$ is much smaller. In the present calculations, we set the energy range of interest to $E_{\gamma}^{\textrm{max}}=50$ MeV. A phenomenological adjustment comes in the enhancement factor assumed in the present calculations, as $\kappa=0.14$ for $sd$-shell nuclei [21] and $\kappa=0$ for $p$-shell nuclei.
@@ III.1 $p$-shell nuclei
III.2 $sd$-shell nuclei
Figure 2: Same as Fig. 1 for $sd$-shell nuclei
@@ III.2 $sd$-shell nuclei
In Ref. [21], we presented systematic calculations of PSFs in long-lived $sd$-shell nuclei and compared them to available data and D1M+QRPA predictions from Ref. [18]. In this work we extended the CI-SM calculations to provide predictions for all 141 existing isotopes that can be described in this valence space. Figs. 2 and 3 compare our CI-SM results for $sd$-shell nuclei close to the stability line with those obtained with other models, namely D1M+QRPA, SMLO [15, 16] and RQFAMz [49]. It can be observed that, while in certain cases the CI-SM predictions are in reasonable agreement with the SMLO and RQFAMz PSFs, in many others the CI-SM results exhibit a higher degree of fragmentation and slight differences in strength magnitude, width, and centroid, which may influence photodisintegration cross sections. Fig. 3 shows the evolution of the centroid and the width as a function of $A$. The decrease of the centroid energy with increasing mass is consistently observed across all three microscopic models with a similar trend. It is worth noting that both D1M+QRPA and RQFAMz include global phenomenological adjustments affecting not only the centroid position but also the folding width. The CI-SM centroids exhibit less smooth variations compared to linear-response-based approaches, as they can incorporate rapid changes in the shell structure and in the GS correlations. Similarly, the CI-SM widths are more scattered and show no clear systematic trend, whereas the D1M+QRPA and RQFAMz widths exhibit a smooth mass dependence. In Ref. [50], we presented a systematic comparison of CI-SM centroids and widths with experimental data. The CI-SM centroid predictions appeared slightly lower than the measured values, with a root-mean-square (rms) deviation of $0.85$ MeV across the 26 cases (see Table 1). For the CI-SM widths, the corresponding rms deviation is $0.50$ MeV, with no systematic under- or over-estimate [^1] Note that the centroid and the width as defined in Eq. 5 are independent of the normalization of the PSF strength described in Sec. II.2.. In CI-SM, we account for the Landau damping as well as most of the spreading width; therefore, the calculated width is almost not sensitive to the folding, which only accounts for the escape width. For comparison, the rms deviation in D1M+QRPA predictions amounts to $1.71$ MeV on the centroids and $0.49$ MeV on the widths, while RQFAMz leads to an rms deviations of $0.89$ MeV on the centroids and $0.48$ MeV on the widths. Hence, even without empirical adjustments to the centroids or widths, the CI-SM predictions remain slightly more accurate than RQFAMz PSFs for $sd$-shell nuclei. One may note that in many cases the experimental PSFs themselves are mutually inconsistent (see Fig. 2), so these rms values should be interpreted with caution. We shall also recall that the shell-model PSDPF effective interaction [28] was only adjusted to reproduce the low-energy spectroscopic data. Therefore, a possible improvement in the PSF predictions could be to incorporate in the fit of the interaction also the measured centroids and widths. From a purely practical perspective one could apply a simple centroid shift, in analogy with the procedure used in QFAM/QRPA calculations. Because the uncorrected CI-SM reproduces the known centroids better than QRPA-based models with empirical corrections, we chose to base our predictions for nuclei used in Sec. IV on the bare CI-SM results.
@@
Figure 3: Panel (a): centroids (up to $30$ MeV) of $E1$ distributions as a function of the mass number AA for long-lived $sd$-shell nuclei. Panel (b): width (up to $30$ MeV) of $E1$ PSFs as a function of the mass number AA for long-lived $sd$-shell nuclei. In red CI-SM, in orange D1M+QRPA [31], in blue RQFAMz results [49].
@@ III.3 O and Ne isotopic chains
After comparing nuclei close to the stability line, we turn to the predictions in the neutron-rich region, where nuclear models tend to show the largest discrepancies. In Figs. 4 and 5, we compare the PSF predictions from all four models for oxygen and neon isotopic chains. As observed earlier, the CI-SM PSFs exhibit significantly more fragmentation than the linear-response-based approaches. While D1M+QRPA and RQFAMz show a smooth evolution of the PSFs with increasing neutron number, the CI-SM dipole responses display sharper variations. Clearly, the SMLO predictions were not tailored for neutron-rich nuclei and therefore cannot account for any PDR. In all three microscopic models, we observe an accumulation of low-lying dipole strength with increasing neutron number, characteristic of the emergence of a PDR. Although it may seem at first sight that no PDR appears in the D1M+QRPA case, it is, in fact, dissolved in the GDR tail. In Refs. [51, 50], we showed that the CI-SM prediction for the PDR in 26Ne is very close to the experimental values (both strength and centroid), which is promising for extending such predictions to other neutron-rich nuclei. Fig. 6 shows the centroids (calculated up to $30$ MeV) for all four models. As can be seen, D1M+QRPA, RQFAMz, and SMLO exhibit the same trend as a function of neutron excess, displaying a gradual decrease. The difference in absolute values between D1M+QRPA and RQFAMz can be attributed to the fact that RQFAMz predicts significantly more strength at high energies (20203030 MeV) than D1M+QRPA. Additionally, phenomenological corrections made in D1M+QRPA PSFs optimize the centroid energies in medium and heavy mass nuclei but tend to underestimate them for light nuclei [18]. Interestingly, the CI-SM centroids follow a very different trend from the other models, exhibiting a much steeper decrease with increasing $A$. This effect is also evident in Fig. 7, where we show the fraction of the PDR EWSR calculated up to $15$ MeV relative to the TRK sum rule [^2] The choice of $15$ MeV as the upper bound, rather than $10$ MeV, is motivated by the fact that the experimental values in Ref. [52] are reported up to this energy. As a consequence, this strength is not purely PDR since it incorporates part of the GDR tail.. Clearly, our calculations predict a stronger increase in the low-lying strength as a function of neutron excess. For the oxygen isotopic chain, experimental data on the PDR are available for a few isotopes [52]. As can be seen in the left panel of Fig. 7, the CI-SM predictions agree quite well with the experiment up to 20O. While they deviate for 21O and 22O, the overall trend remains consistent with the data. This deviation can be attributed to the systematic underestimation of the PSF centroid discussed in the previous section. Indeed, introducing a $1$ MeV upward shift would lead to a much better agreement with the data. While the differences in nuclear models observed for neutron-rich nuclei appear substantial, they have no impact on the modeling of the UHECR, which is dominated by the description of stable and proton-rich nuclei (see Sec. IV). Similar model-to-model variations also appear on the proton-rich side, although the proximity of the proton drip line suppresses the most divergent cases.